# Efficient and Accurate Spatial Mixing of Machine Learned Interatomic Potentials for Materials Science

### v1

@@ 2.2 Energy barrier for Si vacancy migration
A simple zero temperature (static) quantity of interest for a materials defect is the energy barrier for migration. Here, we aimed to show that it is possible to attain a near identical energy barrier for vacancy migration in a $10\times 10\times 10$ (8000 atom) block of Si for both a fully-expensive simulation and an ML/ML simulation where the expensive potential is only localised around the vacancy. For both simulations, the minimum energy pathway was found using the nudged elastic band (NEB) method [49], as implemented in LAMMPS, with further analysis in ASE [50]. For the ML/ML NEB, the combined relaxation was performed using ML-MIX, with the final energies obtained through one-shot energy evaluations of the relaxed structures with the expensive potential. Fig. 3 shows the results from this investigation. The energy barrier obtained through the ML/ML simulation (0.0548 eV) agrees with the all-expensive reference energy barrier (0.0543 eV) to within 1 meV. The NEB relaxation ran $\sim$5$\times$ faster, close to the theoretical ‘zero-overhead’ value of 5.90×5.90$\times$5.90 × expected from the fraction of expensive atoms in the ML/ML simulation.
@@ 2.3 Stretched bond in silicon
Figure 4: Average force on stretched bond in silicon over 100 ps at 300 K for different expensive potential radii $r_{\mathrm{core}}$ and thermostat damping strength. Error bars represent the standard deviation in bond force measured over this time. For $R_{\mathrm{core}}=0$, which corresponds to using the cheap potential everywhere, the average bond force measured does not match the all-expensive $NVE$ reference. Once the expensive potential is introduced at and around the stretched bond, this difference vanishes and agreement between the mixed simulation and the $NVE$ reference is within statistical error. The weaker the thermostat, the better the agreement with the $NVE$ reference.
Turning to a dynamical example, we investigated the quantity of interest proposed as a test for the quality of different QM/MM techniques by Bernstein et al [35] — the average force on a single stretched and rigidly fixed Si bond in a block of bulk Si at 300 K.
In a $10\times 10\times 10$ (8000 atom) supercell of bulk Si, a bond was stretched an additional 0.1 Å along its length and fixed. The average force on this stretched bond recorded over 100 ps was then found, both for all-expensive reference and a number of ML/ML simulations. In the ML/ML simulations, the impact of changing the radius of the expensive potential region around the stretched bond atoms $r_{\mathrm{core}}$ on the measured average force was investigated. The first simulation, at $r_{\mathrm{core}}=0$ corresponded to an all-cheap simulation. Further simulations were conducted for $r_{\mathrm{core}}=4$, $r_{\mathrm{core}}=6$, $r_{\mathrm{core}}=8$ and $r_{\mathrm{core}}=10$.
@@
The theoretical and measured speedups are shown in table 4. It can be seen that in serial, the measured speedups match the theoretical, whilst in parallel the values are lower due to imperfect load balancing and parallel overheads. As $r_{\mathrm{core}}$ increases, the number of expensive atoms in the simulation increases, and the overall speedup decreases. The load balancing strategy employed in this study, as well the effectiveness of load balancing for ML/ML simulations across different regimes is discussed in appendix B.
Table 4: Speedup table for mixed simulations in Si.
@@ 2.4 Diffusion coefficients
A key property of point defects in materials are diffusion coefficients, and estimating these is a common use-case for MD simulations. Here, we consider the diffusion of a dumbbell self-interstitial in Fe and an interstitial He impurity in W, both of which have been extensively studied [51, 52, 53, 54].
For both target defects, reference simulations were carried out using the expensive ACE potentials. These were followed by ML/ML simulations where the expensive potentials were limited to a small region around the diffusing defect, which was tracked as it moved, with the rest of the simulation domain treated using the Fe or W cheap potentials.
Fig. 5 presents the results from these investigations side-by-side. For the Fe system, measurements of the diffusion coefficient as a function of temperature were taken with only the cheap potential, only the expensive potential and using the ML/ML approach. For W-He, only the ML/ML and expensive only simulation results are presented as as the UF3 potential was fit only for pure W and could not model W-He interactions. For both quantities of interest, the ML/ML simulations agreed well with the all-expensive reference calculations at multiple temperatures, despite the inaccuracy of the all-cheap simulation in the Fe case. Further details on the simulation setup are given in section 4.6. The theoretical and measured speedups of the mixed simulations over the reference for both simulation types are shown in table 5. The measured serial speedups again closely agree with the theoretical values whilst the parallel speedups are lower.
Figure 5: Diffusion coefficient as a function of temperature for the diffusion of the iron interstitial dumbbell defect through bulk iron and for helium in bulk tungsten. Left: The diffusion coefficient of the iron dumbbell for three cases: expensive reference potential only (blue), cheap potential only (orange) and mixed simulation (green). The mixed simulation leads to results which are well matched with the reference. Each point is the average of 5 diffusion coefficients measured in independent runs of 1 ns each. The error-bars on each point reflect the standard deviation of these values. Right: The diffusion coefficient for He in W using a helium-tungsten ACE (blue) and a mixed simulation with both a helium-tungsten ACE and a UF3 tungsten potential. Again, agreement is good. Note that there is no ’cheap-only’ reference in this case because the UF3 potential cannot be used to model tungsten-helium interactions. Each point is the average of 25 diffusion coefficients measured in independent runs of 60 ps each. The error-bars on each point reflect the standard deviation of these values.
Table 5: Speedup table for mixed simulations of diffusion coefficients in Fe and W–He.
@@ 3 Discussion
We have presented a hybrid simulation scheme which aims to break existing cost-accuracy tradeoffs in simulations by localizing accuracy requirements, enabling a 10×10$\times$10 × speed-up on small (8000 atom) domains with minimal loss in accuracy on quantities of interest. A central innovation is the use of a constrained fitting procedure to automatically generate cheap models which locally approximate the potential energy surface of a reference expensive model. We show that even for very lightweight ACE potentials, imposing hard constraints can lead to exact matching of elastic constants. For UF3 potentials, we show that under hard constraints, elastic constants will improve up to some limit, beyond which any further improvements are limited by the inflexibility, or misspecification, of the functional form. We stress that whilst we have specifically discussed constrained fitting in the context of generating one MLIP that locally approximates another, it would be easy to apply this methodology to constrain the fitting of a large, flexible MLIP to DFT data directly, extending recent retraining approaches [38].
@@
Constrained fitting is important for this process; without it, errors are far more pronounced at small strains. This can lead to erroneous Eshelby-inclusion[55] like strain fields in the vicinity of the ML/ML potential boundary under small deformations. This is explored in more detail in appendix C.
@@
The first case-study we present in this paper was a nudged elastic band energy barrier calculation for the migration of an Si vacancy. We showed that, using an ML/ML simulation, a value for the migration barrier can be attained which is within 1 meV of the all-expensive reference at a fifth of the computational cost. This speedup, whilst already considerable, is not as great as what is seen for ML/ML MD simulations on the same 8000 atom domain – this is because a larger fixed set of expensive atoms is necessary for a NEB to converge. It is important to point out that total energies are undefined in ML-MIX ML/ML simulations; the final energies were computed by performing one-shot energy evaluations of the relaxed structures using only the expensive potential. In principle, one could take this a step further and perform a small number of additional relaxation steps using only the expensive potential. If the domain was so large or the expensive potential so expensive that even the final single-point energy calculations became intractable, a further option to approximate energy barriers is to directly integrate the ML/ML forces on the final NEB image states using the virtual work method introduced by Swinburne and Kermode [34], originally devised for QM/MM simulations.
@@
Next, we presented an investigation into the average force on a stretched and rigidly fixed bond in an Si bulk structure. This was selected to mirror previous tests performed in QM/MM studies [35]. It should be noted that for an $r_{\mathrm{core}}$ of 0.0 (cheap potential only) the average force measured on the stretched bond is within 5% of the correct value, which is a significantly smaller error than one typically sees in QM/MM simulations. Adding even a small amount of expensive potential region around the stretched bond led to an immediate jump in the measured average bond force, leading to a value which was within $1$ standard deviation ($\sigma$) of the $NVE$ reference value. This behavior persisted for larger $r_{\mathrm{core}}$ values.
The measured serial speed-up closely matched theoretical predictions, showing that our implementation incurs minimal overhead due to the mixing process.
However, in parallel testing (on 48 cores) the measured speedup is below theoretical predictions, in particular for the $r_{\mathrm{core}}=4$ case. We attribute this performance to the limited strong scalability of the expensive interatomic potential at low atom count, leading to significant load balancing issues in LAMMPS when there are few expensive atoms. The limited speed-up is thus a property of the expensive interatomic potential rather than our load balancing protocol- near-theoretical parallel speed-ups are achievable on larger systems with simple load balancing requirements. For a demonstration of this, please see appendix B.
@@
We also presented case studies on more challenging quantities of interest; defect diffusion coefficients in Fe and W. The measured diffusion coefficients consistently agreed with the all-expensive reference simulations within statistical error, and were attained for a fraction of the cost of the full expensive simulations. The close matching of the measured diffusion coefficients with the reference values suggest that one only requires spatially local accuracy to attain correct defect diffusion dynamics in materials; corroborating findings from earlier QM/MM studies [56].
Through these studies, we have demonstrated the flexibility and efficiency of the ML-MIX package to reliably and automatically track defects both in simple cases where the same atoms are always involved in the defects (i.e, for He in W) and for more complex cases where different atoms are involved as the defects move (i.e, for Fe dumbbell interstitial in Fe). Whilst the presented results only concern ACE and UF3 potentials, the general method is extendable to many other interatomic potentials implemented as LAMMPS pair styles. A full list of those compatible can be found on the ML-MIX GitHub repository [57].
@@
It is crucial however to point out a key limitation in ML/ML (and indeed, QM/MM) simulations — if a force mixing method is used to combine potentials (as it is here), Newton’s third law is broken at the internal ML/ML boundary. This directly results in simulations not conserving energy, meaning thermostats are required to keep simulations from heating up as they progress. The energy drift in well-matched ML/ML simulations is relatively small, meaning thermostats can be weak enough to not impact dynamics — but this should always be checked. Force mixing is discussed further in section 4.3.1, and a full investigation into the factors affecting energy conservation in ML/ML simulations is presented in appendix D.
In summary, the ML-MIX approach presented here provides a largely automated and robust scheme to reduce the cost of any machine learned potential with minimal loss in accuracy. We anticipate the approach will find application in a broad range of atomic simulations of heterogeneous systems, including the calculation of defect free energy barriers [58], interface migration [59] or nucleation phenomena [60].

### v2

@@ 2.2 Energy barrier for Si vacancy migration
We first selected a simple zero temperature (static) quantity of interest for a materials defect. We measured the energy barrier for vacancy migration in a $10\times 10\times 10$ (8000 atom) block of Si for both a fully-expensive simulation and an ML/ML simulation where the expensive potential was localised around the vacancy. For the ML/ML NEB, the combined relaxation was performed using ML-MIX. Fig. 2 shows the results from this investigation. The energy barrier obtained through the ML/ML simulation (0.0548 eV) agrees with the all-expensive reference energy barrier (0.0543 eV) to within 1 meV. The NEB relaxation ran $\sim 5\times$ faster, close to the theoretical ‘zero-overhead’ value of $5.90\times$ expected from the fraction of expensive atoms in the ML/ML simulation.
@@ 2.3 Stretched bond in silicon
Figure 3: Average force on stretched bond in silicon over 100 ps at 300 K for different expensive potential radii $r_{\mathrm{core}}$ (blue points). Error bars represent the standard error in bond force measured over this time, with samples taken every 15 fs. For $r_{\mathrm{core}}=0$, which corresponds to using the cheap potential everywhere, the average bond force measured does not match the all-expensive $NVE$ reference (red dashed line). Once the expensive potential is introduced at and around the stretched bond, this difference vanishes and agreement between the mixed simulation and the $NVE$ reference is within statistical error.
Turning to a dynamical example, we investigated the quantity of interest proposed as a test for the quality of different QM/MM techniques by Bernstein et al [Bernstein_2009] — the average force on a single stretched and rigidly fixed Si bond in a block of bulk Si at 300 K.
In a $10\times 10\times 10$ (8000 atom) supercell of bulk Si, a bond was stretched an additional 0.1 Å along its length and fixed. The average force on this stretched bond recorded over 100 ps was then found, both for all-expensive reference and a number of ML/ML simulations, where the impact of changing the radius of the expensive potential region around the stretched bond atoms $r_{\mathrm{core}}$ on the measured average force was investigated. The first simulation, at $r_{\mathrm{core}}=0$ Å, corresponded to an all-cheap simulation. Further simulations were conducted for $r_{\mathrm{core}}=4$ Å, $r_{\mathrm{core}}=6$ Å, $r_{\mathrm{core}}=8$ Å and $r_{\mathrm{core}}=10$ Å. A core region size of 2 Å was not tested as a core radius smaller than the Si-Si bond length of 2.35 Å results in no expensive atoms, and thus an identical result to the $r_{\mathrm{core}}=0$ Å case.
@@
The theoretical and measured speedups are shown in Table 4. It can be seen that in serial, the measured speedups match the theoretical, whilst in parallel the values are lower due to imperfect load balancing and parallel overheads. As $r_{\mathrm{core}}$ increases, the number of expensive atoms in the simulation increases, and the overall speedup decreases. Load balancing is further discussed in section S2 of the supplementary information.
Table 4: Speedup table for mixed simulations in Si. The measured serial speedup closely tracks the theoretically predicted value, whilst the parallel speedup is worse than expected for small $r_{core}$ due to imperfect load balancing.
@@ 2.4 Diffusion coefficients
Next, we moved to a less artificial dynamical quantity of interest. Here, we consider the diffusion of a dumbbell self-interstitial in Fe and an interstitial He impurity in W, both of which have been extensively studied [Wang_2012, Hammond_2018, Willaime_2005, Marinica_2011]. For each of these quantities, an atomistic domain of ($\sim 8000$ atoms) was used. Whilst domains of this size are not necessary for accurate quantities of interest in these simple cells (we found one can attain converged diffusion coefficients in cells of $\sim 500$ atoms and smaller, depending on the defect), we selected this size primarily for demonstration purposes; it is around the minimum cell size beyond which we believe ML/ML methods are worth implementing.
For both target defects, reference simulations were carried out using the expensive ACE potentials. These were followed by ML/ML simulations where the expensive potentials were limited to a small region around the diffusing defect, which was tracked as it moved. Fig. 4 presents the results from these investigations side-by-side. For the Fe system, measurements of the diffusion coefficient as a function of temperature were taken with only the cheap potential, only the expensive potential and using the ML/ML approach. For W-He, only the ML/ML and expensive-only simulation results are presented, as the UF3 potential was fit only for pure W and could not model W-He interactions. For both quantities of interest, the ML/ML simulations agreed well with the all-expensive reference calculations at multiple temperatures, despite the inaccuracy of the all-cheap simulation in the Fe case. The theoretical and measured speedups of the mixed simulations over the reference for both simulation types are shown in Table 5.
Figure 4: Diffusion coefficient as a function of temperature for the diffusion of the iron interstitial dumbbell defect through bulk iron and for helium in bulk tungsten. (a): The diffusion coefficient of the iron dumbbell for three cases: expensive reference potential only (blue), cheap potential only (orange) and mixed simulation (green). The mixed simulation leads to results which are well matched with the reference. Each point is the average of 5 diffusion coefficients measured in independent runs of 1 ns each. (b): The diffusion coefficient for He in W using a helium-tungsten ACE (blue) and a mixed simulation with both a helium-tungsten ACE and a UF3 tungsten potential (orange). Again, agreement is good. Note that there is no ’cheap-only’ reference in this case because the UF3 potential cannot be used to model tungsten-helium interactions. Each point is the average of 25 diffusion coefficients measured in independent runs of 60 ps each. The error-bars on the points in both (a) and (b) represent the standard error in each value.
Table 5: Speedup table for mixed simulations of diffusion coefficients in Fe and W–He. The measured serial speedup closely tracks the theoretically predicted value, whilst the parallel speedup is lower due to imperfect load balancing.
@@
2.5 Velocity of screw dislocation in W
Figure 5: Simulation set-up and results for thermally activated glide of screw dislocations in W at different shear stresses and temperatures. (a) Simulation set-up, where dimensions in $x$, $y$ and $z$ are $223~\mathrm{\AA }\times 221~\mathrm{\AA }\times 68~\mathrm{\AA }\ (25~\textit{b})$, respectively, for a total of 195,000 atoms. Stress was applied to the free surfaces at the top and bottom of the cell in $\langle 111\rangle$ type directions. The expensive atoms around the dislocation (as detected using common neighbor analysis in LAMMPS) are shown in a column, where the colors correspond to the core (red), blending (green) and buffer (white) regions. (b) The velocity results at a range of temperatures and shear stresses. Green, orange and blue lines represent 300, 600 and 900 K respectively. Results from all-expensive simulations are shown with dashed lines, whilst those from mixed simulations are shown with solid lines. For each point, three separate simulations were conducted and the results were averaged. Results from the mixed simulations agree well with those obtained from the all-expensive reference. The error-bars on the points represent the standard error in each value.
The thermally activated glide of screw dislocations is known to play an important role in the deformation of W, with molecular dynamics simulations being a key tool used to understand their mechanics [Bacon_2009]. Following Bacon et al [Bacon_2009] we have used a periodic array of dislocations with Burgers vector $\it{b}$ = $\frac{1}{2}\langle 111\rangle$ to study the thermally activated dislocation glide velocities at a range of applied shear stresses (between 600 and 2000 MPa) and temperatures (300, 600 and 900 K). In each case, all-expensive simulations were conducted using the 3, 21 W ACE potential [Nutter_2024] at a limited range of stresses before ML/ML simulations were conducted using ML-MIX for a greater range, with the expensive region covering only the dislocation.
The simulation domain is presented in Fig. 5(a). The ML-MIX expensive region is highlighted. Fig. 5(b) presents the measured velocity-stress profiles obtained in both the all-expensive and mixed simulations. Each data point is the mean of an ensemble of three repeats, with the error bars representing the standard error. The mixed and all-expensive curves agree with each other within error. The all-expensive simulations in this case were approximately $4\times$ the cost of the mixed simulations.
2.6 He implantation into W
Tungsten is a well-known candidate for plasma-facing materials within modern fusion reactors. Proposed as both a diverter and first-wall material [Abernethy_2017, Khripunov_2015], one of its key properties is the ability to withstand high-energy flux from ions and neutrons. Part of this flux includes impaction from He, the fusion product. In order to predict the damage level and subsequently the longevity of the W material, it is crucial to know (i) how much of the He is expected to enter the W for different incident energies and (ii) how deep the He penetrates. In a previous molecular dynamics study, it was found that these quantities are highly dependent on the choice of interatomic potential [Borovikov_2014], and so far no reconciliation has been made between molecular dynamics results and the limited experimental data that exists [Borovikov_2014, vanGorkum_1980].
To compare with experiment, we specifically targeted simulations of He implantation into {100} W surfaces at 1000 K. Initially, we performed all-expensive simulations using a fine-tuned version of the MACE-mpa-0 foundation model [Batatia_2024]. We then conducted ML/ML simulations using in which the expensive potential interactions were limited to the immediate vicinity of the He atom. In both all-expensive and ML/ML simulations, all computation was conducted on GPU. Fig. 6 shows the measured reflection coefficient as a function of incident He energy. The all-expensive (mpa-0-medium-finetuned) data (blue points), ML-MIX data (orange points) and experimental data (green points) all broadly agree within error. Average implantation depths were also measured for each energy. These are shown in Fig. 7, where it can be seen that results from the ML/ML simulation are in broadly good agreement with the all-expensive. The reduced cost of the ML/ML simulations allowed extension to larger system sizes and thus greater He energies and implantation depths.
[…]
Figure 6: Reflection coefficient as a function of incident He energy in eV for normal deposition onto a 1000 K W {100} W surface. Experimental results from Van Gorkum et al are shown in green [vanGorkum_1980]. Results from the all-expensive potential (fine-tuned mpa-0-medium model) are displayed in blue, and results obtained with ML-MIX are shown in orange. Results from a previous MD simulation using a W EAM and W-He pair potential are displayed in purple [Borovikov_2014]. For the all expensive, to limit computational cost, only energies up to 40 eV were tested with 100 repeats each. With ML-MIX, energies up to 100 eV were repeated 500 times each. Error bars are obtained from the 95% confidence interval of a fitted Beta distribution.
Figure 7: Implantation depth as a function of incident He energy in eV for normal deposition onto a 1000 K W {100} W surface. Results from the all-expensive potential (fine-tuned mpa-0-medium model) are displayed in blue, and results obtained with ML-MIX are shown in orange. For the all expensive, to limit computational cost, only energies up to 40 eV were tested with a 100 repeats each. With ML-MIX, energies up to 100 eV were repeated 500 times each. Error bars represent the standard error in the mean values.
@@ 3 Discussion
We have presented ML-MIX, a hybrid simulation package which aims to break existing cost-accuracy tradeoffs in simulations by localizing accuracy requirements, enabling simulation speed-ups of $4-10\times$ with minimal loss in accuracy on quantities of interest. ML-MIX has the following advantages that set it apart from previously published similar ML/MM schemes: (i) ML-MIX is an easy to install, direct plugin to LAMMPS, which is fast and has no external dependencies, and aims to be compatible with a large number of existing pair_styles. (ii) ML-MIX functions as an external wrapper around LAMMPS pair_styles, meaning there is no necessary editing of C++ code, (iii) ML-MIX works with existing LAMMPS parallelism and load-balancing, and (iv) ML-MIX (though our Kokkos implementation) is compatible with pair_styles that execute on both CPU and GPU.
@@
We have described two ways in which cheap potentials can be generated; either by fitting directly to the same underlying DFT data as the expensive reference, or by fitting to synthetic data generated by the expensive potential in a targeted region of the potential energy surface. We found that fitting to synthetic data was particularly important for very lightweight potentials, as for these we observed that attempting to fit to the diverse reference DFT data directly lead to unacceptably high RMSEs. To ensure close matching of specific properties, we developed a constrained fitting scheme, the details of which are described in section 4.1. For linear ACE potentials, we found that imposing constraints allowed for exact matching of elastic constants between the cheap and expensive potentials. For UF3 potentials, a less flexible architecture, we have shown that under hard constraint elastic constants improve up to some limit, beyond which any further improvements were limited by the misspecification of the functional form. We stress that whilst in this study we have specifically discussed constrained fitting in the context of distilling one MLIP from another, it would be easy to apply this methodology to constrain the fitting of a large, flexible MLIP to DFT data directly, extending recent retraining approaches [Grigorev_2023].
@@
We found that constrained fitting is particularly important for small potentials of <50<50 basis functions; without it, errors are far more pronounced at small strains. This can lead to erroneous Eshelby-inclusion [eshelby1957] like strain fields in the vicinity of the ML/ML potential boundary under small deformations. This is explored in more detail in section S3 of the supplementary information.
@@
The first case-study we present in this paper was a nudged elastic band energy barrier calculation for the migration of an Si vacancy. We showed that, using an ML/ML simulation, a value for the migration barrier can be attained which is within 1 meV of the all-expensive reference at a fifth of the computational cost. This speedup, whilst already considerable, is not as great as what is seen for ML/ML MD simulations on the same 8000 atom domain – this is because a larger fixed set of expensive atoms is necessary for a NEB to converge. It is important to point out that total energies are undefined in ML-MIX ML/ML simulations; the final energies were computed by performing one-shot energy evaluations of the relaxed structures using only the expensive potential.
@@
Next, we presented an investigation into the average force on a stretched and rigidly fixed bond in an Si bulk structure. This was selected to mirror previous tests performed in QM/MM studies [Bernstein_2009]. It should be noted that for an $r_{\mathrm{core}}$ of 0 Å (cheap potential only) the average force measured on the stretched bond is within 5% of the correct value, which is a significantly smaller error than one typically sees in QM/MM simulations. Adding even a small amount of expensive potential region around the stretched bond led to an immediate improvement in the measured average bond force, leading to a value which was within $1$ standard error of the $NVE$ reference value. This behavior persisted for larger $r_{\mathrm{core}}$ values. The measured serial speed-up closely matched theoretical predictions, showing that our implementation incurs minimal overhead due to the mixing process. However, in parallel testing (on 48 cores) the measured speedup is below theoretical predictions, in particular for the $r_{\mathrm{core}}=4\ \mathrm{\AA }$ case. We attribute this performance to the limited strong scalability of the expensive interatomic potential at low atom count, leading to significant load balancing issues in LAMMPS when there are few expensive atoms. The limited speed-up is thus a property of the expensive interatomic potential rather than our load balancing protocol; near-theoretical parallel speed-ups are achievable on larger systems with simple load balancing requirements. This is demonstrated in section S2 of the supplementary information.
@@
We also presented case studies on defect diffusion coefficients in Fe and W. The measured diffusion coefficients consistently agreed with the all-expensive reference simulations within statistical error. The close matching of the measured diffusion coefficients with the reference values suggest that one only requires spatially local accuracy to attain correct defect diffusion dynamics in materials; corroborating findings from earlier QM/MM studies [Csányi_2004].
Next, we considered some more complex and scientifically relevant case-studies, aiming to demonstrate the type of simulations where we believe ML-MIX will find application. These were (i) a study of the thermal glide of screw dislocations in W at a range of temperatures and pressures, and (ii) a study of normal deposition of He in W at 1000 K.
We considered the thermally activated glide of $\it{b}$ = $\frac{1}{2}\langle 111\rangle$ screw dislocations in W. We found that in a mixed ACE/ ACE simulation, we were able to reproduce the dislocation velocities measured in the all-expensive reference, meaning that the mixed simulation was able to reproduce the correct dynamics of the kink-nucleation-propagation mechanism that controls the glide of screw dislocations in W. Whilst there was scope to study longer dislocation lines, we found increasing the length of the dislocation increased the probability of cross-kinking – a phenomenon that can lead to debris formation. This can cause nonphysical interference with subsequent dislocation motion when the dislocation re-encounters the debris due to periodicity. As such, we leave an exploration of larger dislocation studies with ML-MIX to future work.
Finally, we simulated the normal implantation of He into (100) W surfaces at He energies up to 70 eV at 1000 K. This led to a novel result; for the first time, experimental observations of reflection coefficient were reproduced up to 40 eV with the mpa-0-medium foundation model fine-tuned to W–He DFT data. Following this, we carried out the same simulations with ML-MIX, where we limited interactions with the expensive potential to directly around the He and used a pure W SNAP potential for the remainder of the interactions. In order to ensure efficient evaluation of the expensive MACE potential, ML-MIX was ported to GPU using Kokkos [kokkos1, kokkos2]. Using ML-MIX, we showed that (i) up to 40 eV results agreed well with the all expensive simulation and experiment, and (ii) above 40 eV results continue to line up closely with experiment, with deviations only occuring at relatively high energy (>80>80 eV). Measurements of implantation depth were also taken in each case, and those from the mixed simulations aligned well with those from the all-expensive simulations. Due to the relative low cost of the mixed simulations, we were able to perform more repeats and create the larger cells necessary for high energy simulations without needing to parallelize over more GPUs.
It is important to note that the experiment we have drawn comparisons with involved incident He+ ions [vanGorkum_1980], whilst in this study we are considering incident He atoms. One possible explanation for our close agreement despite this is that in reality, as the He+ approaches the surface, it gains an electron from the W conduction band on timescales faster than the movement of the nuclei. A recent time-dependent DFT study by Ito et al [Ito_2025] studying plasma-wall interactions observed electron transfer in the case of He2+, but was inconclusive on He+. We believe that reconciling this and performing a wider exploration of the implantation parameter space are both beyond the scope of this primarily method-focused paper, and therefore leave them as interesting directions for future work.
@@
A key limitation in ML/ML (and indeed, QM/MM) simulations arises from the lack of energy conservation; if a force mixing method is used to combine potentials (as it is here), simulations do not conserve energy. This means that a thermostat may be required to keep simulations from substantially heating up as they progress. The energy drift in well-matched ML/ML simulations is relatively small ($\mathcal{O}(10~\mathrm{K/ps})$), meaning thermostats can be weak enough to not impact dynamics — but this should always be checked. Force mixing is discussed further in section 4.5, and a full investigation into the factors affecting energy conservation in ML/ML simulations is presented in section S4 of the supplementary information.
A possible solution to this problem is the use of so-called ‘energy-mixing’ [Bernstein_2009], where a Hamiltonian is constructed for the mixed system, for example by writing the total energy as a sum of local site energies from each potential. Upon testing this scheme however, we found the forces derived had errors of >50>50% in the vicinity of the ML/ML boundary, even when using potentials that nominally are well matched in force-mixing. This backs up previous work on energy-mixing for QM/MM conducted by Bernstein et al [Bernstein_2009] and expanded on by Chen & Ortner [Chen_Ortner_2017] where it was found that eliminating these spurious ‘ghost forces’ in energy-mixed QM/MM for infinitesimal displacements away from equilibrium required the MM potential have an exactly matching force-constant matrix (second order derivatives of the site energies with respect to atomic positions). This requirement is stringent and as such would require more flexible (and thus expensive) cheap potentials, and even if satisfied only would eliminate ghost forces for atoms near equilibrium. Additionally, if the regions change over time as atoms move (a common feature of long-timescale ML/ML simulations), energy would not be conserved anyway; as regions update, potential energy would be continually added or removed from atoms. Conversely, a buffered force-mixing approach naturally has no spurious ghost-forces and can be made to work with far simpler cheap potentials. In view of this, we believe (despite the lack of energy conservation) that force-mixing is the clear choice for ML/ML simulations.
A further limitation is that we have made no attempt to quantify or estimate how the error associated with the size of the expensive region selected may propagate through to quantities of interest. Uncertainty quantification (UQ) for MLIPs and QM/MM is a very active area of research [Swinburne_2025, Perez_2025, Best_2024, Wang_2024], and we believe that applications of UQ to ML/ML simulations is a very exciting possible direction for future work.
Through these studies, we have demonstrated the flexibility and efficiency of the ML-MIX package. In our initial, simple case-studies, we show that ML-MIX can be used to accelerate simulations without compromising accuracy both in simple cases where the same atoms are always involved in the defects (i.e, for He in W) and for more complex cases where different atoms are involved as the defects move (i.e, for Fe dumbbell interstitials in Fe). In two further case-studies, we have shown that ML-MIX can be used to attain a significant speedup without any meaningful loss in accuracy in applications of genuine scientific interest. In the first of these, we present results for the speed of $\it{b}$ = $\frac{1}{2}\langle 111\rangle$ screw dislocations in W at a range of temperatures and shear stresses. In the second, we provide estimates of reflection coefficients and implantation depths for normally deposited He into 1000 K {100} W surfaces. Our measured reflection coefficients match experimental observations up to a deposition energy of 80 eV. Looking forwards, we anticipate that ML-MIX will find application in a broad range of atomic simulations of heterogeneous systems, including the calculation of defect free energy barriers [Swinburne_2018], interface migration [Rogal_2021] and nucleation phenomena [Bonati_2021].